# Adoption of Watermarking Measures for AI-Generated content and Implications under the EU AI Act

### v2

@@ 4.3. Detection of watermarks and disclosures
To analyse the adoption of marking and disclosure measures in practice (in the context of the AI Act), we used the two evaluation metrics outlined in Table 1.
Table 1. Metrics used for evaluating the implementation of marking and labelling measures in AI systems
@@ 5. Results
Figure 4 shows an overview of the machine-readable marking solutions that were found in our analysis. The most used solution was the embedding of metadata that indicates the AI-generated origin of the data (12 systems), which is easily removable from the images. In five of these instances, the embedded metadata was part of the C2PA standard. Four out of the five providers using this standard also embedded a digital fingerprint in the image, so that the C2PA verify tool can be used to inspect the provenance of the image. Invisible watermarking techniques were found in only a small subset of the systems (8 systems). Two of these belonged to the end-to-end system providers Google and StabilityAI, as also indicated in their documentation. The other six instances of invisble watermarks were not disclosed by the providers but were found using the open-source python watermarking library that is used in the popular open-source models of Stability AI and Black Forest Labs. However, in many of the other systems that rely on the models (or APIs) from Stability AI or Black Forest Labs, we could not detect this watermark.
@@
Visible markings for deep fake images are rarely used in practice. Visible watermarks or other disclosure solutions that are embedded within the image and that indicate the AI-generated nature of the content were found in only 9 out of the 50 systems (18%). Figure 5 shows three examples of such disclosures in images generated with the neutral prompt. In all nine cases, these markings were not exclusively applied to deep fake images, but rather to all the images generated with the system. In five out of the nine systems that used visible markings, there was an option for end-users to remove the visible markings, either through a checkbox or via signing-up to a paid plan, as illustrated in Figure 6.
Interestingly, while end-to-end providers (category 1) have almost all implemented a form of machine-readable marking, visible disclosures (of deep fake images) were found to be used in only 2 of the 10 end-to-end systems. Some of these end-to-end systems, however, did restrict the use of clear deep fake prompts (e.g., Donald Trump), showing an alternative approach to tackling the risks of deep fakes.
@@
Figure 7. Overview of the main countries of listing of the system providers and the different GPAI models that were indicated to be used in the systems.
@@ 6. Limitations
It is essential to acknowledge several limitations and constraints in our analysis. First, there is a possibility that we failed to identify certain watermarking and digital fingerprinting techniques that are used in practice. We tried our best to find any disclosures of watermarking from the documentation of the providers of the systems, but there is a chance that they might have not disclosed them and used non-standard techniques that we did not check for. Second, this study has a temporal limitation. The generative AI systems and machine-readable marking solutions that we examined in this study are regularly updated, and the implementation of marking and labelling practices is likely to change over time. Third, although we used prompts that we clearly see as deep fakes, we admit that the definition of a deep fake might be subject to debate. Finally, some AI systems that we analysed did not offer a save or download button for the images within the systems’ interface environment, which required us to use other downloading methods. This may have caused certain metadata to be lost, potentially affecting our findings regarding the embedding of machine-readable watermarks.
@@ 7. Discussion: The wild west of AI image generation
Markings and labels are rarely used in practice. Our findings show that only a minority of providers currently implement machine-readable marking practices, as will soon be required by the EU AI Act. It is predominantly large and well-funded organisations (mostly from category 1) that implement such solutions. Given that many of these organisations also operate digital and social media platforms, the implementation of machine-readable marking measures may primarily be driven by aligning incentives, such as preventing the pollution of their (social) media platforms or AI training datasets, or protecting IP rights. Visible disclosures indicating the AI-generated nature of images are even more rarely used, especially among end-to-end operators from category 1. This is likely driven by the impact that such measures have on the user experience, and not because of costs or technical difficulties. At the same time, we see that several smaller organisations do use visible markings to get users to sign-up to a paid plan that allows the removal of the markings (which is unlikely to satisfy AI Act compliance). When visible disclosures are used, they are applied to all generated images, rather than specifically deep fake images, as soon mandated by the AI Act.
Lack of robust machine-readable marking, even among large providers. Given that the AI Act requires robust machine-readable marking solutions (as far as technically feasible), it is questionable if most of the currently used techniques comply with this requirement. This is particularly the case for providers that exclusively embed metadata (6 systems), which can be easily removed. Robust watermarking methods that focus on detecting AI generated images, remain rare, even among large providers. Additionally, the majority of marking solutions are currently relying on post-generation techniques. Using robust watermarking methods directly in the model generation phase, particularly by category 1 GPAI model providers, could significantly ease implementation for downstream system providers using APIs and require significant extra effort for bad actors to remove. The absence of such pre-embedded solutions now forces smaller system providers to either implement cheaper and weaker methods (metadata), or to refrain from implementing any solutions at all, as observed in many cases in this study.
Concentrated AI ecosystem, with just a few players. Our findings highlight the important role that providers of the most advanced (open-source) models play in shaping the ecosystem. A large number of system providers rely on models (or fine-tuned versions) from just a handful of providers (e.g., Stability AI, Black Forest Labs and OpenAI). If the providers of these widely used image models can design more robust watermarking solutions that are not easily bypassed, this could shift the paradigm, requiring model deployers to actively remove watermarks, rather than add them. Platforms such as Hugging Face can play an important role in enforcing such measures for open-source models that are hosted on their platform. Especially as Hugging Face might also qualify as a deployer (and possibly provider) of generative AI systems when developers use the API inference widget to deploy the model on the Hugging Face platform. By not exempting open-source systems from the transparency requirements, the AI Act has thus taken an important step to address the critical role of open-source providers.
@@
Challenges in restricting labelling to deep fakes, especially for smaller organisations. To restrict the use of visible disclosures to specifically deep fake images, instead of all AI-generated images, providers would need a separate system - likely NLP-based - that can classify certain prompts as deep fakes, which can prove difficult and computationally expensive for smaller organisation to implement. We do, however, observe that similar methods are used in practice, as several large organisations restrict specific deep fake prompts (e.g., Donald Trump), showing an alternative method to tackle the harms of deep fakes. Whether visible markings will become more widely implemented as the AI Act’s transparency requirements take effect in August 2026 remains uncertain, especially considering that most of the providers of popular generative AI systems are not based in the EU. Moreover, it is necessary to consider the question of how effective some of the current markings (as seen in Figure 4) are in helping the normal layperson understand that these images are AI-generated.
@@ 8. Conclusions and how to move beyond the status quo
In this paper, first, we studied the implications of the EU AI Act on the need to implement watermarking techniques for AI-generated images and deep fakes. We identified four categories of AI system that are relevant under the AI Act and outlined the legal obligations for each category. Next, we studied the status quo in implementing watermarking in practice, looking at 50 of the most widely used AI systems for image generation. We found that and only few providers currently implement robust watermarking solutions and that visible AI deepfake labels are rarely used.
@@
Our legal analysis highlights the ambiguities around the distinctions made in the AI Act between providers, deployers, models and systems in the context of the generative AI image systems. The burden for compliance with the machine-readable marking and deep fake disclosure requirements may disproportionally fall on small-scale app and website developers, rather than on the large-scale and well-funded organisations that develop the models behind these applications. Yet, robust machine-readable solutions would be best implemented at the model development stage, where they could also be enforced via API and licensing terms. Plus, enforcing compliance at the model developer level would be easier given the relatively small number of organisations that can develop advanced image models. One potential solution could be to designate the most advanced image models as GPAI models with systemic risk, which would require developers to take measures for mitigating systemic risks, such as the use of robust watermarks. Alternatively, it could be considered to use the implementing guidelines of the AI Act to classify providers that offer ready-to-use model APIs as providers of AI systems, as it could be argued that such an API is a way of adding an interface to a model and creating a system. This would require model API providers to implement robust machine-readable marking solutions in the model stage.
@@
Implementing the AI Act’s visible disclosure requirement for deep fakes poses additional challenges, which could lead to risky large-scale non-compliance gaps. Restricting disclosures to only deep fake images requires complex methods that could be hard to implement for smaller developers, but applying labels to all generated images could negatively impact the user experience and draw users to platforms of non-compliant providers. Finally, given the variety of marking techniques that are used in practice, and the growing number of deployed AI image generation systems (often deployed from non-EU countries), automated methods for compliance inspection will be essential to ensure compliance and effective enforcement when the AI Act’s requirements come into effect.
@@
Moving forward, society is faced with an uphill battle where it will be challenging to have adequate watermarks in all AI-generated images. After all, much of the relevant AI technology is available open-source and free of charge. With models getting smaller, it will increasingly become easier to produce AI-generated images at home. However, it is reassuring that the largest providers generative AI systems do already provide protections. It will, however, be extremely difficult to protect society from actors with malicious intents in challenging our perceptions of truth and reality. In addition, while this study focused on images, the AI Act also requires the implementation of adequate marking solutions in video, audio and text generation systems, for which the implementation challenges could prove even harder.

### v3

@@ 4.3. Detection of watermarks and disclosures
To analyse the adoption of marking and disclosure measures in practice (in the context of the deployment categories defined in our legal analysis), we used the two evaluation metrics outlined in Table 1.
Table 1. Metrics used for evaluating the use of marking and labelling measures in AI systems
@@ 5. Results
Figure 4 shows an overview of the machine-readable marking solutions that were found in our analysis. The most used solution was the embedding of metadata that indicates the AI-generated origin of the data (12 systems), which is easily removable from the images. In five of these instances, the embedded metadata was part of the C2PA standard. Four out of the five providers using this standard also embedded a digital fingerprint in the image, so that the C2PA verify tool can be used to inspect the provenance of the image. Invisible watermarking techniques were found in only a small subset of the systems (8 systems). Two of these belonged to the end-to-end system providers Google and StabilityAI, as also indicated in their documentation. The other six instances of invisible watermarks were not disclosed by the providers but were found using the open-source python watermarking library that is used in the popular open-source models of Stability AI and Black Forest Labs. However, in many of the other systems that rely on the models (or APIs) from Stability AI or Black Forest Labs, we could not detect this watermark.
@@
Visible markings for deep fake images are rarely used in practice. Visible watermarks or other disclosure measures that are embedded within the image or AI system and that can help indicate the AI-generated nature of the content were found in only 9 out of the 50 systems (18%). Figure 5 shows three examples of such visible disclosures in images generated with the neutral prompt. In all nine cases, these markings were not exclusively applied to deep fake images, but rather to all the images generated with the system. In five out of the nine systems that used visible markings, there was an option for end-users to either add the visible marking through a checkbox, or remove the markings by signing-up to a paid plan, as illustrated in Figure 6.
Interestingly, while end-to-end providers (category 1) have almost all implemented a form of machine-readable marking, visible disclosures were found to be used in only 2 of the 10 end-to-end systems. Some of these end-to-end systems, however, did restrict the use of clear deep fake prompts (e.g., Donald Trump), showing an alternative approach to tackling the risks of deep fakes.
@@
Figure 7. Overview of the main countries of listing of the system providers (left) and the different GPAI models that were indicated to be used within the systems (right).
@@ 6. Limitations
It is essential to acknowledge several limitations and constraints in our empirical analysis. First, there is a possibility that we failed to identify certain watermarking and digital fingerprinting techniques that are used in practice. We tried our best to find any disclosures of watermarking from the documentation of the providers of the systems, but there is a chance that they might have not disclosed them and used non-standard techniques that we did not check for. Second, this study has a temporal limitation. The generative AI systems and machine-readable marking solutions that we examined in this study are regularly updated, and the implementation of marking and labelling practices is likely to change over time. Third, although we used prompts that we clearly see as deep fakes, we admit that the definition of a deep fake might be subject to debate. Finally, some AI systems that we analysed did not offer a save or download button for the images within the systems’ interface environment, which required us to use other downloading methods. This may have caused certain metadata to be lost, potentially affecting our findings regarding the embedding of machine-readable watermarks.
@@ 7. Discussion: The wild west of AI image generation
Markings and labels are rarely used in practice. Our findings show that only a minority of providers currently implement machine-readable marking practices, as will soon be mandated by the EU AI Act. It is predominantly large and well-funded organisations (mostly from category 1) that implement such solutions. Given that many of these organisations also operate digital and social media platforms, the implementation of machine-readable marking measures may primarily be driven by aligning incentives, such as preventing the pollution of their (social) media platforms or AI training datasets, or protecting IP rights.
@@
Visible disclosures indicating the AI-generated nature of images are even more rarely used, especially among end-to-end operators from category 1. This is likely driven by the impact that such measures have on the user experience, and not because of costs or technical difficulties. At the same time, we see that several smaller organisations do use visible markings to get users to sign-up to a paid plan that allows the removal of the markings. When visible disclosures are used, they are applied to all generated images, rather than specifically deep fake images. To restrict the use of visible disclosures to specifically deep fake images, instead of all AI-generated images, providers would need a separate system - likely NLP-based - that can classify certain prompts as deep fakes, which can prove difficult and computationally expensive for smaller organisation to implement. We do, however, observe that similar methods are used in practice, as several large organisations restrict specific deep fake prompts (e.g., Donald Trump). Whether visible markings will become more widely implemented as the AI Act’s transparency requirements take effect in August 2026 remains uncertain, especially considering that most of the providers of popular generative AI systems are not based in the EU. Moreover, it is necessary to also consider the question of how effective some of the current visible markings (as seen in Figure 5) are in helping the normal layperson understand that these images are AI-generated.
Concentrated AI ecosystem, with just a few players. Our findings highlight the important role that providers of the most advanced (open-source) models play in shaping the ecosystem. A large number of system providers rely on models (or fine-tuned versions) from just a handful of GPAI model providers (e.g., Stability AI, Black Forest Labs and OpenAI). These providers do try to incorporate robust watermarking solutions in their models, and we find several instances of such solutions in systems relying on these models. However, these solutions are also quite easy to disable (e.g., by commenting out a line of code) and not consistently applied to API-based deployment methods. If these providers can design more robust watermarking solutions that are not easily bypassed, this could shift the paradigm, requiring model deployers to actively remove watermarks, rather than add them. Platforms such as Hugging Face can play an important role in enforcing such measures for open-source models that are hosted on their platform. Especially as Hugging Face might also qualify as a provider (and possibly deployer) of generative AI systems, when the Hugging Face widgets are used to deploy a model on the Hugging Face platform. By not exempting open-source systems from the transparency requirements, the AI Act has thus taken an important step in to address the critical role of open-source providers.
Lack of robust machine-readable marking, even among large providers. Given that the AI Act requires robust machine-readable marking solutions (as far as technically feasible), it is questionable if most of the currently used techniques comply with this requirement. This is particularly the case for providers that exclusively embed metadata (6 systems), which can be easily removed. Robust watermarking methods that focus on detecting AI generated images, remain rare, even among large providers. Additionally, the majority of machine-readable marking solutions are currently relying on post-generation techniques. Using robust marking methods directly in the model generation phase, particularly by category 1 GPAI model providers, could significantly ease implementation for downstream system providers using APIs and require significant extra effort for bad actors to remove. The absence of such pre-embedded solutions now forces smaller system providers to either implement cheaper and weaker methods (metadata), or to refrain from implementing any solutions at all, as observed in many cases in this study.
@@ 8. Conclusions and how to move beyond the status quo
In this paper, first, we studied the implications of the EU AI Act on the need to implement watermarking techniques for AI-generated images and deep fakes. We identified four categories of AI system that are relevant under the AI Act and outlined how the legal obligations could apply for each category. Next, we studied the status quo in implementing watermarking in practice, looking at 50 of the most widely used AI systems for image generation. We found that and only few providers currently implement robust watermarking solutions and that visible AI deepfake labels are yet rarely used in practice.
Our legal analysis highlights the ambiguities around the distinctions made in the AI Act between providers, deployers, models and systems in the context of the generative AI image systems, and the considerable impact that further guidance on these distinctions can have. One example is the interpretation of the term deployer in the context of the deep fake disclosure requirement. Deployer can either refer to (professional) end-users of a system, or to entities ‘deploying’ a system to such end-users. Here, it would be most practical and impactful to implement disclosure measures at the system level (such as giving users an option to add a label), rather than having end-users manually search and apply such labels themselves after generation.
@@
Similarly, the burden for compliance with the machine-readable marking measure may disproportionally fall on small-scale app and website providers, rather than on the large-scale and well-funded organisations that develop the models behind these applications. Yet, robust marking and labelling solutions would be best implemented at the model development stage, where they could also be enforced via API and licensing terms. Plus, enforcing compliance at the model developer level would be easier given the relatively small number of organisations that can develop advanced image models. One potential solution could be to designate the most advanced image models as GPAI models with systemic risk, which would require developers to take measures for mitigating systemic risks, such as the use of robust marking measures for deep fake images. Alternatively, it could be considered to use the implementing guidelines of the AI Act to classify providers that offer ready-to-use model APIs as providers of AI systems, as it could be argued that such an API is a way of adding an interface to a model and creating a system. This would require model API providers to implement robust machine-readable marking solutions in the model stage.
@@
Finally, given the variety of marking techniques that are used in practice, and the growing number of deployed AI image generation systems (often deployed from non-EU countries), automated methods for compliance inspection will be essential to ensure compliance and effective enforcement when the AI Act’s requirements come into effect.
@@
Moving forward, society is faced with an uphill battle where it will be challenging to have adequate watermarks in all AI-generated images. After all, much of the relevant AI technology is available open-source and free of charge. With models getting smaller, it will increasingly become easier to produce AI-generated images at home. However, it is reassuring that the largest providers generative AI systems do already provide protections. It will, however, be extremely difficult to protect society from actors with malicious intents in challenging our perceptions of truth and reality. In addition, while this study focused on images, the AI Act also requires the implementation of adequate marking and labelling solutions in video, audio and text generation systems, for which the implementation challenges could prove even harder.